# Mehdi Zafari A. Lee Swindlehurst

### v1

@@ Abstract
Integrated Sensing and Communication (ISAC) is a key emerging 6G technology. Despite progress, ISAC still lacks scalable methods for joint AP clustering and user/target scheduling in distributed deployments under fronthaul limits. Moreover, existing ISAC solutions largely rely on centralized processing and full channel state information, limiting scalability. This paper addresses joint access point (AP) clustering, user and target scheduling, and AP mode selection in distributed cell-free ISAC systems operating with constrained fronthaul capacity. We formulate the problem as a mixed-integer linear program (MILP) that jointly captures interference coupling, RF-chain limits, and sensing requirements, providing optimal but computationally demanding solutions. To enable real-time and scalable operation, we propose ASSENT (ASSociation and ENTity selection), a graph neural network (GNN) framework trained on MILP solutions to efficiently learn association and mode-selection policies directly from lightweight link statistics. Simulations show that ASSENT achieves near-optimal utility while accurately learning the underlying associations. Additionally, its single forward pass inference reduces decision latency compared to optimization-based methods. An open-source Python/PyTorch implementation with full datasets is provided to facilitate reproducible and extensible research in cell-free ISAC.
@@ I Introduction
Integrated Sensing and Communication (ISAC) is emerging as a core paradigm for next-generation wireless systems, simultaneously serving communication users (CUs) and environmental sensing needs using shared spectrum and hardware resources [liu2022dual-functional]. Cell-free architectures [ngo2017cellfree], which eliminate predefined cell boundaries and enable cooperation among access points (APs), was originally proposed for communication networks to enhance spectral efficiency and manage interference more effectively. This architecture is now being extended to ISAC systems [umut2025cellfreeisac], where it offers additional advantages for sensing, such as larger effective apertures, improved clutter mitigation, and higher parameter estimation accuracy.
@@
Several recent studies have leveraged distributed optimization techniques to address resource allocation in cell-free networks [zafari2024admm, zafari2025decentralized]. However, a key bottleneck in extending these architectures to ISAC lies in the absence of efficient association and entity selection mechanisms that can jointly determine which APs serve which CUs and targets, and in which operational modes. The scheduling problem for ISAC systems has been investigated in a few recent works assuming full network knowledge [time-domain-scheduling-2024, setareh2025scheduling]. In parallel, graph learning-based methods have emerged as promising tools for modeling cooperative ISAC networks [peng2025graph, wang2024GNN]. Nevertheless, existing approaches typically assume centralized computation with global information, treat communication and sensing tasks in isolation, overlook spatial interference among CUs, or often rely on heuristic strategies that lack adaptability to dynamic and large-scale network deployments.
In this work, we consider a distributed cell-free ISAC network with constrained fronthaul capacity. To limit the fronthaul load, each AP reports low-dimensional statistics for its communication and sensing links to a CS. These lightweight statistics inform an association problem that consists of joint AP clustering, user and target scheduling, and AP mode selection, which is formulated as a mixed-integer linear program (MILP). To capture network dynamics and enable real-time decision-making, we propose ASSENT (ASSociation and ENTity selection), a graph neural network (GNN)-based framework trained on offline solutions of the MILP that learns the desired association behavior. Once trained, ASSENT performs fast forward inference to deliver real-time decisions, shifting the heavy optimization burden offline and avoiding per-slot MILP implementation. We also provide an open-source Python/PyTorch implementation of the proposed framework, together with the training and evaluation datasets. The contributions of this work are summarized below:
@@ II System and Channel Model
We consider a cell-free network consisting of a set $\mathcal{A}$ of $N_{ap}$ distributed APs, a set $\mathcal{U}$ of $N_{cu}$ CUs, and a set $\mathcal{T}$ of $N_{tg}$ targets. Each AP $a$ is equipped with an antenna array of size $M_{a}$, with a limited number of RF chains $N_{\mathrm{RF}}^{(a)}$, where $N_{\mathrm{RF}}^{(a)}\leq M_{a}$. The system employs orthogonal frequency-division multiplexing (OFDM) and operates in time-division duplex (TDD) mode. A resource block (RB) is defined as a contiguous portion of time-frequency resources consisting of $Q_{f}$ subcarriers in the frequency domain and $Q_{t}$ OFDM symbols in the time domain. We consider a downlink scenario in which a subset $\mathcal{A}_{t}$ of APs transmits data to CUs and illuminates targets for sensing, while another subset $\mathcal{A}_{r}$ of APs operates in receive mode to capture the reflected target echo signals. We consider half-duplex operation mode for the APs, i.e., $\mathcal{A}_{t}\cap\mathcal{A}_{r}=\emptyset$, to avoid self-interference.
@@ II-A Communication Channel Model
To model the communication channels, we adopt a narrowband stochastic channel based on the 3GPP Urban Micro (UMi) street canyon model [3gpp_r18]. For each user-AP link, a distance-dependent line-of-sight (LoS) probability is defined according to the UMi scenario. The LoS state determines the corresponding path loss, and statistical parameters such as the mean and standard deviation (STD) of the angle spreads, shadow fading, and the Rician K-factor are drawn following the specifications in [3gpp_r18]. A correlated shadowing model is employed to capture spatial coupling among CUs. The covariance between the shadow fading variables of CUs $j$ and $k$ as seen from AP $a$ is expressed as $\text{cov}(\psi_{a,j},\psi_{a,k})=\sigma_{a,j}\sigma_{a,k}\exp(-d_{j,k}/D_{\text{corr}})$, where $\psi_{a,j}$ denotes the shadow fading random variable, $\sigma_{a,j}$ its STD, $d_{j,k}$ the inter-user distance, and $D_{\text{corr}}=50\,\mathrm{m}$ the decorrelation distance.
@@
$\rho_{u,u^{\prime}}^{a}[p,q]=\frac{|\widehat{\mathbf{h}}_{a,u}[p,q]^{H}\,\widehat{\mathbf{h}}_{a,u^{\prime}}[p,q]|^{2}}{\|\widehat{\mathbf{h}}_{a,u}[p,q]\|^{2}\cdot\|\widehat{\mathbf{h}}_{a,u^{\prime}}[p,q]\|^{2}},$ (2)
@@ II-B Sensing Channel Model
For the sensing links, we assume the same UMi street canyon environment as for UEs. Target $t$ is located at distance $d_{a,t}$ from AP $a$. The LoS probability between $a$ and $t$ is determined using the 3GPP distance-dependent LoS model. To characterize the target radar cross section (RCS), we adopt the statistical measurements reported in [nextGalliance_tr_phase2]. Each AP $a$ reports to the CS the following parameters for each target $t$: 3D AP-target distance $d_{a,t}$, estimated RCS $\bar{\beta}_{a,t}^{\text{rcs}}$ averaged over a given RB, and estimated LoS state, inferred from either the distance or prior measurements. At the CS, for each bistatic sensing link formed by a transmit AP $a_{t}$, a target $t$, and a receive AP $a_{r}$, an approximate channel gain is computed as
$\bar{g}_{a_{t},t,a_{r}}=\$$\frac{G_{\text{tx}}G_{\text{rx}}\lambda_{w}^{2}}{(4\pi)^{3}}\cdot$ (3)
$\frac{1}{d_{a_{t},t}^{2}d_{t,a_{r}}^{2}}\left(\frac{\bar{\beta}_{a_{t},t}^{\text{rcs}}+\bar{\beta}_{t,a_{r}}^{\text{rcs}}}{2}\right)W(p_{a_{t},t}^{\text{LoS}})W(p_{t,a_{r}}^{\text{LoS}}),$
where $G_{\text{tx}}$ and $G_{\text{rx}}$ denote the antenna gains of the transmitting and receiving APs, respectively, and $\lambda_{w}$ is the carrier wavelength. These three parameters are assumed to be known at the CS. The term $W(p^{\text{LoS}})$ is a weighting function that modulates the contribution of each link based on its LoS probability:
@@ III-A Optimization Variables
The association between AP $a$ and user $u$ is represented by a binary variable $x_{a,u}\in\{0,1\}$, where $x_{a,u}=1$ indicates that AP $a$ serves user $u$ in the downlink, and $x_{a,u}=0$ otherwise. For target $t$, a binary association variable $y_{a,t}^{\text{tx}}\in\{0,1\}$ is defined to indicate whether AP $a$ is transmitting to illuminate target $t$ ($y_{a,t}^{\text{tx}}=1$) or not, and a variable $y_{a,t}^{\text{rx}}\in\{0,1\}$ is defined to indicate whether AP $a$ is receiving echo signals from target $t$ ($y_{a,t}^{\text{rx}}=1$) or not. The operation mode of AP $a$ is also determined by variable $\tau_{a}\in\{0,1\}$, where $\tau_{a}=1$ indicates transmit mode and $\tau_{a}=0$ receive mode. Additionally, a binary variable $s_{t}\in\{0,1\}$ is introduced to determine whether target $t$ is scheduled ($s_{t}=1$) or not. If scheduled, target $t$ must be associated with two non-overlapping subsets of APs, $\mathcal{A}_{t}$ for illumination and $\mathcal{A}_{r}$ for echo reception.
@@ III-B Objective Function
The association problem objective is to maximize a weighted sum of utility functions for CUs and targets. For each $k\in\mathcal{U}\cup\mathcal{T}$, representing either a CU or a target, a weight $\lambda_{k}$ is assigned based on its service requirements and historical scheduling frequency, ensuring fairness and preventing any entity from being persistently under-served. The communication utility function is accordingly defined as
@@ IV-A Dataset Generation
To train the proposed ASSENT model, we generate a comprehensive dataset of association and mode-selection decisions by solving the MILP in (7) over a large number of independent channel realizations. For each realization, the system parameters summarized in Table I include the communication channel gains $\mathbf{G}_{\text{comm}}$, inter-user spatial correlations $\mathbf{S}_{\text{comm}}$, and the bistatic sensing link gains $\mathbf{G}_{\text{sens}}$, as well as the control parameters $\alpha$, $\boldsymbol{\lambda}_{\text{cu}}$, and $\boldsymbol{\lambda}_{\text{tg}}$. The MILP is solved centrally at the CS to obtain the globally optimal binary decision variables $\{\tau,\,x,\,s,\,y^{\text{tx}},\,y^{\text{rx}}\}$. Each realization therefore provides a complete input-output pair used for supervised learning.
@@
$\mathbf{G}_{\text{comm}}$ Comm. channel gains (AP-User) $\mathbb{R}^{N_{ap}\times N_{cu}}$
$\mathbf{S}_{\text{comm}}$ User-User spatial correlations $\mathbb{R}^{N_{ap}\times N_{cu}\times N_{cu}}$
@@
$\boldsymbol{\lambda}_{\text{cu}}$ Vector of user priority weights $\mathbb{R}^{N_{cu}}$
@@
$\mathbf{X}$ AP-User association $\{0,1\}^{N_{ap}\times N_{cu}}$
@@
To generate channel realizations, we consider a network with $N_{ap}=8$ APs, $N_{cu}=10$ users, and $N_{tg}=4$ targets. The APs are uniformly placed on a circle of radius $350$ m, while users and targets are randomly positioned within a $1$ km $\times$ 11 km area centered around the APs. Each AP is equipped with $N_{\text{RF}}=4$ RF chains and a UCA with $M=16$ antennas. The carrier frequency is $3$ GHz, and the trade-off parameter $\alpha$ is randomly drawn from $[0,1]$. Targets are modeled as cars with RCS statistics drawn from [nextGalliance_tr_phase2] and velocities uniformly distributed in $[0,10]$ m/s. The target RCS is modeled as a class-dependent random variable with mean value $\bar{\beta}_{\text{class}}\in[10,20]$ dBsm. To capture fluctuations due to the aspect angle between the target and the APs, a random variable $\Delta_{\text{asp}}\sim\mathcal{N}(0,\,\sigma_{\text{asp}}^{2}(v))$ is introduced, whose STD $\sigma_{\text{asp}}(v)$ varies with the target velocity $v$ as
@@
where $v_{\text{ref}}$ is the reference velocity sensitivity (set to $1\,\text{m/s}$) and $v_{\text{max}}$ is the maximum velocity for the target class (set to $30\,\text{m/s}$). The tunable parameters $\sigma_{\text{asp}}^{\text{min}}$ and $\sigma_{\text{asp}}^{\text{max}}$ control the impact of velocity on RCS fluctuations. Accordingly, each realization of the RCS for the AP-target link is expressed as $\beta_{a,t}^{\text{rcs}}\,[\text{dB}]=\bar{\beta}_{\text{class}}+\Delta_{\text{asp}}$. The priority weights $\lambda_{u}$ and $\lambda_{t}$ for each user $u$ and target $t$ are uniformly sampled from $\lambda_{\min},\lambda^{\max}]$ for $\lambda_{\min}=0.5$ and $\lambda^{\max}=\{5,10\}$. Using these parameters, a total of 10,000 independent realizations are generated by randomizing user and target positions, propagation conditions, and control parameters. The dataset is divided into training, validation, and test subsets with an 80, 10, 10 ratio. Each realization is serialized as a heterogeneous graph object and stored in binary format for efficient loading during training. This dataset captures a statistically representative set of scenarios, enabling the learning model to generalize to unseen network configurations and trade-off conditions.

### v2

@@ Abstract
Integrated sensing and communication (ISAC) is a key emerging 6G technology. Despite progress, ISAC still lacks scalable methods for joint access point (AP) clustering and user/target scheduling in distributed deployments under fronthaul limits. Moreover, existing ISAC solutions largely rely on centralized processing and full channel state information, limiting scalability. This paper addresses joint AP clustering, user and target scheduling, and AP mode selection in distributed cell-free ISAC systems operating with constrained fronthaul capacity. We formulate the problem as a mixed-integer linear program (MILP) that jointly captures interference coupling, RF chain limits, and sensing requirements, providing optimal but computationally demanding solutions. To enable real-time and scalable operation, we propose ASSENT (ASSociation and ENTity selection), a graph neural network (GNN) framework trained on MILP solutions to efficiently learn association and mode-selection policies directly from lightweight link statistics. Simulations show that ASSENT achieves near-optimal utility while accurately learning the underlying associations. Additionally, its single forward pass inference reduces decision latency compared to optimization-based methods. An open-source Python/PyTorch implementation with full datasets is provided to facilitate reproducible and extensible research in cell-free ISAC.
@@ I Introduction
Integrated sensing and communication (ISAC) is emerging as a core paradigm for next-generation wireless systems, simultaneously serving communication users (CUs) and environmental sensing needs using shared spectrum and hardware resources [5]. Cell-free architectures [8], which eliminate predefined cell boundaries and enable cooperation among access points (APs), were originally proposed for communication networks to enhance spectral efficiency and manage interference more effectively. This architecture is now being extended to ISAC systems [3], where it offers additional advantages for sensing, such as larger effective apertures, improved clutter mitigation, and higher parameter estimation accuracy.
@@
Several recent studies have leveraged distributed optimization techniques to address resource allocation in cell-free networks [11, 10]. However, a key bottleneck in extending these architectures to ISAC lies in the absence of efficient association and entity selection mechanisms that can jointly determine which APs serve which CUs and targets, and in which operational modes. The scheduling problem for ISAC systems has been investigated in a few recent works assuming full network knowledge [6, 2]. In parallel, graph learning-based methods have emerged as promising tools for modeling cooperative ISAC networks [4, 9]. Nevertheless, existing approaches typically assume centralized computation with global information, treat communication and sensing tasks in isolation, overlook spatial interference among CUs, or often rely on heuristic strategies that lack adaptability to dynamic and large-scale network deployments.
In this work, we consider a distributed cell-free ISAC network with constrained fronthaul capacity. To limit the fronthaul load, each AP reports low-dimensional statistics for its communication and sensing links to a CS. These lightweight statistics inform an association problem that consists of joint AP clustering, user and target scheduling, and AP mode selection, which is formulated as a mixed-integer linear program (MILP). To capture network dynamics and enable real-time decision-making, we propose ASSociation and ENTity selection (ASSENT), a graph neural network (GNN)-based framework trained on offline solutions of the MILP that learns the desired association behavior. Once trained, ASSENT performs fast forward inference to deliver real-time decisions, shifting the heavy optimization burden offline and avoiding per-slot MILP implementation. We also provide an open-source Python/PyTorch implementation of the proposed framework, together with the training and evaluation datasets. The contributions of this work are summarized below:
@@ II System and Channel Model
Let us consider a cell-free network consisting of $N_{ap}$ distributed APs indexed by set $\mathcal{A}$, $N_{cu}$ CUs indexed by set $\mathcal{U}$, and $N_{tg}$ targets indexed by set $\mathcal{T}$. Each AP $a$ is equipped with an antenna array of size $M_{a}$, with a limited number of RF chains $N_{\mathrm{RF}}^{(a)}$, where $N_{\mathrm{RF}}^{(a)}\leq M_{a}$. The system employs orthogonal frequency-division multiplexing (OFDM) and operates in time-division duplex (TDD) mode. A resource block (RB) is defined as a contiguous portion of time-frequency resources consisting of $Q_{f}$ subcarriers in the frequency domain and $Q_{t}$ OFDM symbols in the time domain. We consider a downlink scenario in which a subset $\mathcal{A}_{t}$ of APs transmits data to CUs and illuminates targets for sensing, while another subset $\mathcal{A}_{r}$ of APs operates in receive mode to capture the reflected target echo signals. We consider half-duplex operation mode for the APs, i.e., $\mathcal{A}_{t}\cap\mathcal{A}_{r}=\emptyset$, to avoid self-interference.
@@ II-A Communication Channel Model
To model the communication channels, we adopt a narrowband stochastic channel based on the 3GPP Urban Micro (UMi) street canyon model [1]. For each user-AP link, a distance-dependent line-of-sight (LoS) probability is defined according to the UMi scenario, and the LoS state determines the corresponding path-loss. The statistical link parameters such as the mean and standard deviation (STD) of the angle spreads, shadow fading, and the Rician K-factor are drawn following the specifications in [1]. A correlated shadowing model is employed to capture spatial coupling among CUs. The covariance between the shadow fading variables of CUs $j$ and $k$ as seen from AP $a$ is expressed as $\text{cov}(\psi_{a,j},\psi_{a,k})=\sigma_{a,j}\sigma_{a,k}\exp(-d_{j,k}/D_{\text{corr}})$, where $\psi_{a,j}$ denotes the shadow fading random variable, $\sigma_{a,j}$ its STD, $d_{j,k}$ the inter-user distance, and $D_{\text{corr}}=50\,\mathrm{m}$ the decorrelation distance.
@@
$\rho_{u,u^{\prime}}^{a}[p,q]=\frac{|\widehat{\mathbf{h}}^{H}_{a,u}[p,q]\,\widehat{\mathbf{h}}_{a,u^{\prime}}[p,q]|^{2}}{\|\widehat{\mathbf{h}}_{a,u}[p,q]\|^{2}\cdot\|\widehat{\mathbf{h}}_{a,u^{\prime}}[p,q]\|^{2}},$ (2)
@@ II-B Sensing Channel Model
For the sensing links, we assume the same UMi street canyon environment as for CUs. Target $t$ is located at distance $d_{a,t}$ from AP $a$. The LoS probability between $a$ and $t$ is determined using the 3GPP distance-dependent LoS model. To characterize the target radar cross section (RCS), we adopt the statistical measurements reported in [7]. Each AP $a$ reports to the CS the following parameters for each target $t$: 3D AP-target distance $d_{a,t}$, estimated RCS $\bar{\beta}_{a,t}^{\text{rcs}}$ averaged over a given RB, and estimated LoS state, inferred from either the distance or prior measurements. At the CS, for each bistatic sensing link formed by a transmit AP $a_{t}$, a target $t$, and a receive AP $a_{r}$, an approximate channel gain is computed as
$\bar{g}_{a_{t},t,a_{r}}=\$$\frac{G_{\text{tx}}G_{\text{rx}}\,\lambda_{w}^{2}\,{\bar{\beta}}_{a_{t},t,a_{r}}^{\text{rcs}}}{(4\pi)^{3}\,d_{a_{t},t}^{2}\,d_{t,a_{r}}^{2}}\cdot W(p_{a_{t},t}^{\text{LoS}})W(p_{t,a_{r}}^{\text{LoS}}),$ (3)
where $G_{\text{tx}}$ and $G_{\text{rx}}$ denote the antenna gains of the transmitting and receiving APs, respectively, $\lambda_{w}$ is the carrier wavelength, and ${\bar{\beta}}_{a_{t},t,a_{r}}^{\text{rcs}}=(\bar{\beta}_{a_{t},t}^{\text{rcs}}+\bar{\beta}_{t,a_{r}}^{\text{rcs}})/2$ represents the effective RCS obtained by averaging the transmit-target and target-receive links. The gains and the wavelength are known at the CS, and $\bar{\beta}_{a,t}^{\text{rcs}}$, $d_{a,t}$, and $p_{a,t}^{\text{LoS}}$ are provided by APs for all links. The term $W(p^{\text{LoS}})$ is a weighting function that modulates the contribution of each link based on its LoS probability:
@@ III-A Optimization Variables
The association between AP $a$ and CU $u$ is represented by a binary variable $x_{a,u}\in\{0,1\}$, where $x_{a,u}=1$ indicates that AP $a$ serves CU $u$ in the downlink, and $x_{a,u}=0$ otherwise. For target $t$, a binary association variable $y_{a,t}^{\text{tx}}\in\{0,1\}$ is defined to indicate whether AP $a$ is transmitting to illuminate target $t$ ($y_{a,t}^{\text{tx}}=1$) or not, and a variable $y_{a,t}^{\text{rx}}\in\{0,1\}$ is defined to indicate whether AP $a$ is receiving echo signals from target $t$ ($y_{a,t}^{\text{rx}}=1$) or not. The operation mode of AP $a$ is also determined by variable $\tau_{a}\in\{0,1\}$, where $\tau_{a}=1$ indicates transmit mode and $\tau_{a}=0$ receive mode. Additionally, a binary variable $s_{t}\in\{0,1\}$ is introduced to determine whether target $t$ is scheduled ($s_{t}=1$) or not. If scheduled, target $t$ must be associated with two non-overlapping subsets of APs, $\mathcal{A}_{t}$ for illumination and $\mathcal{A}_{r}$ for echo reception.
@@ III-B Objective Function
The objective of the association problem is to maximize a weighted sum of utility functions for CUs and targets. For each $k\in\mathcal{U}\cup\mathcal{T}$, representing either a CU or a target, a weight $\lambda_{k}$ is assigned based on its service requirements and historical scheduling frequency, ensuring fairness and preventing any entity from being persistently under-served. The communication utility function is accordingly defined as
@@ IV-A Dataset Generation
To train the proposed ASSENT model, we generate a comprehensive dataset of association and mode-selection decisions by solving the MILP in (7) over a large number of independent channel realizations. For each realization, the input system parameters summarized in Table I include the communication channel gains $\mathbf{G}_{\text{comm}}$, inter-user spatial correlations $\mathbf{S}_{\text{comm}}$, and the bistatic sensing link gains $\mathbf{G}_{\text{sens}}$, as well as the control parameters $\alpha$, $\boldsymbol{\lambda}_{\text{cu}}$, and $\boldsymbol{\lambda}_{\text{tg}}$. The MILP is solved centrally at the CS to obtain the globally optimal binary decision variables $\{\tau,\,x,\,s,\,y^{\text{tx}},\,y^{\text{rx}}\}$. Each realization therefore provides a complete input-output pair used for supervised learning.
@@
$\mathbf{G}_{\text{comm}}$ Comm. channel gains (AP-CU) $\mathbb{R}^{N_{ap}\times N_{cu}}$
$\mathbf{S}_{\text{comm}}$ CU-CU spatial correlations $\mathbb{R}^{N_{ap}\times N_{cu}\times N_{cu}}$
@@
$\boldsymbol{\lambda}_{\text{cu}}$ Vector of CU priority weights $\mathbb{R}^{N_{cu}}$
@@
$\mathbf{X}$ AP-CU association $\{0,1\}^{N_{ap}\times N_{cu}}$
@@
To generate channel realizations, we consider a network with $N_{ap}=8$ APs, $N_{cu}=10$ CUs, and $N_{tg}=4$ targets. The APs are uniformly placed on a circle of radius $350$ m, while users and targets are randomly positioned within a $1$ km $\times$ 11 km area centered around the APs. Each AP is equipped with $N_{\text{RF}}=4$ RF chains and a UCA with $M=16$ antennas. The carrier frequency is $3$ GHz, and the trade-off parameter $\alpha$ is randomly drawn from $[0,1]$. Targets are modeled as cars with RCS statistics drawn from [7] and velocities uniformly distributed in $[0,10]$ m/s. The target RCS is modeled as a class-dependent random variable with mean value $\bar{\beta}_{\text{class}}\in[10,20]$ dBsm. To capture fluctuations due to the aspect angle between the target and the APs, a random variable $\Delta_{\text{asp}}\sim\mathcal{N}(0,\,\sigma_{\text{asp}}^{2}(v))$ is introduced, whose STD $\sigma_{\text{asp}}(v)$ varies with the target velocity $v$ as
@@
where $v_{\text{ref}}$ is the reference velocity sensitivity (set to $1\,\text{m/s}$) and $v_{\text{max}}$ is the maximum velocity for the target class (set to $30\,\text{m/s}$). The tunable parameters $\sigma_{\text{asp}}^{\text{min}}$ and $\sigma_{\text{asp}}^{\text{max}}$ control the impact of velocity on RCS fluctuations. Accordingly, each realization of the RCS for the AP-target link is expressed as $\beta_{a,t}^{\text{rcs}}\,[\text{dB}]=\bar{\beta}_{\text{class}}+\Delta_{\text{asp}}$. The priority weights $\lambda_{u}$ and $\lambda_{t}$ for each CU $u$ and target $t$ are uniformly sampled from $\lambda_{\min},\lambda^{\max}]$ for $\lambda_{\min}=0.5$ and $\lambda^{\max}=\{5,10\}$. Using these parameters, a total of 10,000 independent realizations are generated by randomizing user and target positions, propagation conditions, and control parameters. The dataset is divided into training, validation, and test subsets with an 80, 10, 10 ratio. Each realization is serialized as a heterogeneous graph object and stored in binary format for efficient loading during training. This dataset captures a statistically representative set of scenarios, enabling the learning model to generalize to unseen network configurations and trade-off conditions.